# Token Spammers, Rug Pulls, and SniperBots: An Analysis of the Ecosystem of Tokens in Ethereum and the Binance Smart Chain (BNB)

### v2

@@ 7.3.2 The Tokens’ Names
To further deepen our analysis of exit scams, we focus on the names used in the frauds. Analyzing the exit scams, we notice several tokens with the same name in BSC and Ethereum. We find that the tokens involved in exit scams have only 157,864 (57.9%) and 18,801 (86.4%) unique names in BSC and Ethereum. Table 4 shows the most used names and the number of occurrences for each of them. Thus, we attempt to cluster the exit scam tokens into categories and enumerate them.
@@
As a first category, we explore clones—tokens with the same name as an existing (and more popular) cryptocurrency. To systematically search for these cases, we need an authoritative source to discern what tokens are legitimate and what are clones. We leverage CoinGecko APIs [16] to retrieve the name and the addresses of all tokens created and verified with the indexer service on the BSC and Ethereum. At the end of the process, we build a list of 5,325 BSC tokens, and 5,172 Ethereum. We complement this list by adding a list of popular variations for some tokens’ names (e.g., , we also considered ADA as a possible name for the Cardano token). Using our list, we discover 22,002 cloned tokens in BSC and 1,781 in Ethereum. The most cloned tokens in BSC are Berryswap (370 occurrences), ApeSwap (210 occurrences), Shiba Inu (191 occurrences), and SafeMoon (158 occurrences).
@@
The second category we explore is the one of tokens that attempt to impersonate companies or websites. In this case, we retrieve the name of the companies of the Standard and Poor’s 500 (S&P 500) stock market index to obtain a list of possible target companies. Instead, for the websites, we extract the second-level domain from the top-ranked 200 websites according to the Alexa ranking  [^5] Data retrieved Apr. 26, 2022. Using in conjunction these two lists, we find 4,638 tokens of this category in BSC and only 95 in Ethereum. The companies and websites that are present the most are Pornhub (1,023), Spacex(419), Onlyfans (398), Oracle (319), and Amazon (270).
@@
We find several repetitions of names that contain popular meme-related words like "Doge", "Inu" or "Shiba". This is not surprising since meme coins are very popular nowadays after events that involved the "meme stocks" of GameStop (GME) and AMC Entertainment (AMC) in late 2020 [34]. Luckily, CoinMarketCap and CoinGecko offer a categorization of the tokens they list that containing also the "meme" category. We leverage these lists to extract the most frequent words and search for them in tokens involved in exit scams. We find a huge amount of tokens in this category: 54,229 in BSC and 4,835 in Ethereum.
@@
As the last category of our investigation, we look for DeFi services (e.g., Deriswap, Shibaswap, Seedswap, and Eco Finance). In this case, we search for tokens containing the "swap", "defi", and "finance" keywords. With this approach, we find for this category 24,197 tokens in BSC and 3,548 in Ethereum.
@@
With our simple categorization, we covered the names of 39% of the exit scam tokens on BSC and 47% in Ethereum. Even if we were not able to categorize all the tokens, we get some insights into how scammers pick the name to arrange their fraud. In particular, it is possible to note a strong trend in choosing tokens’ names related to the meme category and leveraging the names of popular cryptocurrencies, services, and companies.
@@
Figure 4: Each data point represents an address swapping inside liquidity pools with an exit scam pattern. On the y axis, we represent the number of different liquidity pools where the address swaps. On the x axis we show the average time interval between the first time the liquidity is added to the liquidity pool and the swaps operations of the address.
We find that many exit scams are successful without using fake tokens or wash trading. Since these zero-effort operations are very quick and simple, it is still unclear how they can be profitable. We analyze the operations carried out inside exit scams more in-depth and find out that their success may be the activity of a particular class of trading bots, which we analyze in the following.
8.1 Sniper Bots Description
Sniper bots are automated bots that monitor time-bound activities and perform an action before or after anyone else. An example are “Auction Snipers” which are used to place last-second bidding on auction sites such as eBay [47]. In this case, the goal is to secure the highest bid in the auction. Another example of sniper bots is “Scalping Bot”, bots that monitor the availability of target products from a website and buy them as soon as they are available (e.g., GPU Nvidia GPUs) [11]. Using these bots, it is possible to buy a product as soon as it is available and sell it at a later time at a higher price.
With the birth of and the widespread adoption of AMMs, a new kind of sniper bot has been developed that we define Sniper Bots 2.0. These kinds of sniper bots are programs that buy tokens on liquidity pools as soon as they are listed. Thus, the basic idea is to “snipe” a new coin by buying it before anyone else at its initial price. To do so in the fastest way, sniper bots can leverage the mempool— the list of transactions not yet inserted in blockchain blocks.
We find examples of these bots distributed for free on Github [50, 45, 19] and for a price at several other websites [2, 44]. Analyzing the code, we can infer how they work. As a first step, the bot connects to the blockchain network. Then, it continuously scans the mempool looking for transactions whose byte-code indicates that they are adding liquidity to a brand new liquidity pool. The bot sends a swap transaction to buy the token. If the gas price is adjusted correctly, it is executed in the same block (but immediately after) the transaction that adds the liquidity. Sniper bots typically execute only the buy operation. The user can freely decide when to sell the tokens and make a profit. However, we also found some variants that automatically sell the tokens when the price rise by a certain percentage.
@@ 8.2 Identifying Sniper Bots
We conjecture that one of the reasons for the profitability of exit scam operations is sniper bots that buy tokens from every liquidity pool indiscriminately. Thus, we can consider the liquidity pools involved in exit-scam as“honey pots” to detect sniper bots. To verify our intuition, we focus on addresses that have swapped inside liquidity pools with an exit scam. Figure 4 shows the phenomenon: every dot is an address, and its position indicates the number of different liquidity pools where the address has swapped and the average delay from the pool creation. The figure shows a few addresses that swap into thousand of liquidity pools almost immediately after creation. Since these addresses perform these operations serially and incredibly fast, we believe that these addresses must be sniper bots. We set up two conservative thresholds to identify evidence of addresses used by sniper bots. We consider all the addresses that swap on average with a delay smaller than 5 blocks (15 seconds, BSC blocks are mined every 3 seconds on average) and that swap in at least 100 different liquidity pools.
@@
We flag 130 addresses as possible sniper bots. These addresses represent only 0.05% of all the addresses that swap inside liquidity pools involved in exit scams. It is impressive that they swap in 235,777 liquidity pools, representing 86.5% of all the liquidity pools with an exit scam. Moreover, these addresses perform also an impressive number of swaps: 646,339. That account for 24% of all the swaps performed in liquidity pools with an exit scam. We find that 31% of these swaps are performed in the same block where the liquidity is added for the first time in the liquidity pool. In these cases, we can confirm that the sniper bots scanned the mempool to swap in the same block where the liquidity is added. However, we also find sniper bots that perform the swap operations a few blocks after the liquidity is created.
From the code of the open-source sniper bots, we find that many implementations do not leverage the mempool but instead wait for the token to be listed on services like BscScan or Etherscan. Indeed, the documentation of sniper bots that use the mempool reports that it is necessary to deploy an Ethereum or BSC node locally. This requires technical knowledge and hardware that is not accessible to most users.
@@
We find sniper bots to be less present in Ethereum. In this case, we pick two thresholds and consider all the addresses that swap on average with a distance lower than 3 blocks (45 seconds, Ethereum blocks are mined on average every 15 seconds) and that swap in at least 10 liquidity pools. We find 64 possible sniper bots that are only the 0.1% of all Ethereum addresses that swap in exit scam liquidity pools. We find that these addresses swap in 30% of all the liquidity pools and perform a much smaller fraction of swaps than BSC sniper bots (3.5% of the total). However, interestingly, a higher percentage of swaps are performed in the same block where the liquidity is added to the liquidity pools (60%).
@@ 9 Related Work
Tokens identification. In previous work, there are mainly two approaches for token identification: behavior based approach and interface based approach.
In behavior based approach, there is the assumption that a token contract has a data structure that maps addresses to the quantity of token they own and a function to transfer tokens. Chen et al. [14] observe the EVM execution trace to identify in smart contracts data structure that identify the bookkeeping of a token. They also detect inconsistencies between the actual behavior of the token and the actions indicated by standard interfaces, and the behaviors suggested by standard events.
In interface oriented approach, that is the approach that we take in this work, the aim is to identifiy tokens that are compliant with specific interfaces (e.g ERC20 interface). Thus, the approach consist in locating in the smart contract bytecode the functions and events they implement. Several works use this approach [21, 53, 15]. Frowis et al [27] combine both approaches and compare their performances. They demonstrated that they are able to detect 99% of the tokens in their ground truth dataset using the interface based.
ERC-20 tokens graphs. Some studies on ERC-20 tokens focus on graph analysis. Victor et al. [53] introduce the concept of Token Networks to study ERC-20 tokens from a network perspective. Token Networks are graphs where the nodes are addresses that have owned a specific ERC-20 token and they are connected if there are transfers of such tokens between them. They collect a dataset of 75,514 ERC-20 tokens and build a Token network for 64,393 of them. They analyze the top 1000 Token Networks and find out that they account for 85% of all the token transfers and that some of them barely show any activity after the initial token distribution. Chen et al. [15] expand this work by creating transaction graphs of the whole token ecosystem, considering more than 160,000 tokens. Tokens have been explored also on other blockchains, albeit to the best of our knowledge there are no work that focus on the BSC. Zheng et al. [58] analyzed the EOSIO token ecosystem via graph analysis, focusing on the graph of token creators, token contract creators, and token holders. Similarly to other works on Ethereum, they found out that most of the tokens are “silent” and only 1% of the tokens cover more than 90% of the total token volume. They propose a fake-token detection algorithm to detect anomalous manipulation activities, finding three suspect cases.
@@
Liquidity pools frauds. Xia et al. [56] characterize scam tokens on Ethereum, focusing mainly on the Uniswap DEX. First they leverage CoinMarketCap [17] to obtain a ground truth of official and scam tokens. They use The Graph [28] , a sandbox to obtain blockchain data, to obtain 21,778 tokens and 25,131 liquidity pools from May 2020 to December 2020. Firs they expand the dataset by using a guilt-by-association heuristic and then they use it to train a machine-learning model. Then they run their classifier on the extended dataset and find more than 11,182 scam tokens. Mazorra et al. [37] extend the dataset of Xia et al, adding Uniswap data until 3 September 2021, finding other 18 thousand scam tokens.They provide a classification of three different types of rug pulls, simple, sell, and trap-door rug pulls. Then, they found that more than the 97.7% of the tokens labeled as scam are involved in rug pulls operations.
@@ 10 Discussion
What is the impact of not collecting all the internal transactions? Unlike other works [53, 15], we do not collect all smart contracts generated by internal transactions. We collect smart contracts created directly by EOAs, and expand our dataset by adding contracts that emitted at least a Transfer Event. This approach could lead to the loss of a small percentage of tokens. We can perform a rough estimation of the ERC-20 token we miss by comparing the number of tokens we retrieved with the number of tokens retrieved by Chen et al. [15] at the same block height. We find that our approach retrieve 146,928 token instead of 165,955, approximately 12% less. However, it important to note that, by design, our approach misses only tokens that are never used, traded or transferred. So, the missing tokens do not represent interesting cases for our study.
@@
Why does it appear that frauds and token spammers are more frequent in BSC than in Ethereum? From a technical point of view, the frauds work the same way in the two blockchains. Indeed, since BSC is EVM compliant and PancakeSwap is a fork of Uniswap, the same smart contracts can be used on both blockchains. However, the cost of the operation is significantly different. As we saw in Section 7.3.1, performing the fraud in BSC is cheaper (on average \$10.5 with peaks of \$600) than in Ethereum (on average \$400 with peaks of over \$2,000). These costs represent a fixed cost for the scammer, and thus going even or gaining money may be more difficult in Ethereum versus BSC.
@@
Can different users coordinate to carry out the same operation, or can a user use multiple addresses? In this work, we considered each address as belonging to a single different user, and we assumed that there is no coordination among addresses. Nonetheless, a user may occasionally switch or, at regular intervals, the address he uses to perform each fraud. It is also possible that a group of users coordinate to carry out the fraud. For example, a user can create a liquidity pool, and the others perform wash trading. A possible approach to detect this malicious behavior is to gather all the transactions among the allegedly involved address and look for malicious patterns or communities (e.g., using graph analysis). In this work, we do not perform this analysis, but we plan to explore more sophisticated frauds as an extension of this work.
Practical relevance of our work. On the Uniswap and PancakeSwap websites, it is possible to select the tokens to swap using the token’s name, the symbol or the address. In our study, we find plenty of fake tokens that attempt to emulate the original ones. An unaware user that selects the pair using the token’s name or symbol can erroneously select a fake pair. Both protocols already provide some kind of protection to the users, displaying a warning message in case a user selects a not official pair—a pair not created by the protocol itself. While many not official pairs are not a fraud and have legitimate reasons to be traded, we have evidence that token spammers are also serial scammers, and it is better not to trust their tokens. In the light of the results of this work, we believe that swap protocols should implement token spammer detection systems and flag the liquidity pools created by these addresses (or new addresses linked to spammers) to increase the users’ awareness. Sniper bots can build blocklists of untrusted addresses and avoid swapping into liquidity pools created by token spammers.

### v3

@@ 7.2.3 Tokens’ names
To further deepen our analysis of rug pulls, we focus on the names used in the operations. Analyzing the rug pulls, we notice several tokens with the same name in BSC and Ethereum. We find that of the 272,349 tokens involved in the operations in BSC and 21,742 in Ethereum there are only 157,864 (57.9%) and 18,801 (86.4%) unique names. Thus, we attempt to cluster the 1-day tokens into categories and enumerate them. Table 6 in the Appendix shows the most used names and the number of occurrences for each of them.
@@
As a first category, we explore clones—tokens with the same name as an existing (and more popular) cryptocurrency. To systematically search for these cases, we use as an authoritative source the CoinGecko APIs [16]. Leveraging them, we retrieve the names and the addresses of all tokens created and verified with the indexer service on the BSC and Ethereum. At the end of the process, we build a list of 5,325 tokens for BSC, and 5,172 tokens for Ethereum. We complement this list by adding popular variations for some tokens’ names (e.g., we also considered ADA as a possible name for the Cardano token). Using our list, we discover 22,002 cloned tokens in BSC and 1,781 in Ethereum. The most cloned tokens in BSC are Berryswap (370), Shiba Inu (191), and SafeMoon (158).
@@
The second category we explore is the one of tokens that attempt to impersonate companies or websites. In this case, to obtain a list of possible target companies, we retrieve the name of the companies of the Standard and Poor’s 500 (S&P 500) stock market index. Instead, for the websites, we extract from the Alexa ranking  [^7] Data retrieved 2022-04-26 the name of the top-ranked 200 websites. Using in conjunction these two lists, we find 4,638 tokens of this category in BSC and only 95 in Ethereum. The companies and websites that are present the most are Pornhub (1,023), Spacex(419), Onlyfans (398), Oracle (319), and Amazon (270).
@@
We find several names that contain popular meme-related words like "Doge", "Inu" or "Shiba". This is not surprising, since meme tokens are very popular after the events that involved the "meme stocks" of GameStop (GME) and AMC Entertainment (AMC) in late 2020 [45]. Luckily, CoinMarketCap and CoinGecko offer a categorization of the tokens that also contain the "meme" category. We leverage these lists to extract the most frequent words and search for them into the tokens involved in rug pulls. We find a huge amount of tokens of this category: 54,229 in BSC and 4,835 in Ethereum.
@@
As the last category of our investigation, we look for DeFi services (e.g., Deriswap, Shibaswap, and Eco Finance). In this case, we simply search for tokens containing the "swap", "defi" and "finance" keywords. With this approach, we find for this category 25,524 tokens in BSC and 3,751 in Ethereum.
@@
With our simple categorization, we covered the names of 39% of the 1-day tokens on the BSC and 48% on Ethereum. Even if we were not able to categorize all the tokens, we get some insights on how attackers pick the name to arrange their operations. In particular, we note a strong trend in choosing tokens’ names related to the meme category and leveraging the name of popular cryptocurrencies, services, and companies.
@@
Figure 5: Scatter plot of the number of liquidity pools with a 1-day rug pull pattern where the address swapped and the average delay from the pool creation.
We find that a large fraction of rug pulls are successful, even if they are zero-effort operations, without fake tokens or wash trading. Since these kinds of operations are very quick and simple, it is still unclear how they can be profitable. We analyze the operations carried out inside rug pulls more in-depth and discover that their success may be due to the activity of a particular class of trading bots, called Sniper Bots.
Sniper bots are automated bots that monitor time-bound activities and perform an action before or after anyone else. An example of sniper bot are “Scalping Bots”, bots that monitor the availability of target products from a website and buy them as soon as they are available (e.g., Nvidia GPUs) [10].
With the birth of and the widespread adoption of AMMs, a new kind of sniper bot has been developed, which we define Sniper Bots 2.0. These kinds of sniper bots are programs that buy tokens on liquidity pools as soon as they are listed. To do so in the fastest way, sniper bots can leverage the mempool— the list of transactions not yet inserted in blockchain blocks. We find examples of these bots distributed for free on Github [60, 52, 20] and for a price at several other websites [2, 49]. Analyzing the code, we can infer how they work. As a first step, the sniper bot must search for newly listed tokens. The fastest implementation scans the mempool looking for transactions whose byte-code indicates that they are adding liquidity to a brand new liquidity pool. Another possibility is that the sniper bot waits for the token to be listed on services like BscScan or Etherscan. Then, the bot sends a swap transaction to buy the token, and if the gas price is properly adjusted, it is executed in the same block (but immediately after) of the transaction that adds the liquidity. Sniper bots typically execute only the buy operation. The user then can freely decide when to sell the token and make a profit. However, we also found some variants that automatically sell the token when the price reaches a pre-defined goal.
@@ 8.1 Identifying Sniper Bots
We conjecture that one of the reasons for the profitability of rug pulls operations are sniper bots that buy tokens from every liquidity pool indiscriminately. Thus, we can consider the liquidity pools involved in rug pulls as “honey pots” to detect sniper bots. To verify our intuition, we focus on addresses that swapped inside liquidity pools with a rug pull Fig. 5 shows the phenomenon: Every dot is an address, and its position indicates the number of different liquidity pools where the address swapped and the average delay from the pool creation. The figure shows a few addresses that swap in thousand of liquidity pools almost immediately after their creation. Since these addresses perform these operations serially and incredibly fast, we believe they must be sniper bots. We set up two conservative thresholds to identify evidence of addresses used by sniper bots.
@@
For BSC, we consider all the addresses that swap on average with a delay smaller than five blocks (15 seconds) and that swap in at least 100 different liquidity pools. We flag 130 addresses as possible sniper bots. These addresses represent only 0.03% of all the addresses that swap inside liquidity pools involved in rug pulls. What is impressive is that they swap in 235,777 liquidity pools, representing 68.7% of all the liquidity pools with a rug pull. Moreover, these addresses also perform an impressive number of swaps: 2,691,173, that account for 24% of all the swaps performed in liquidity pools with a rug pull. We find that 31% of these swaps are performed in the same block where the liquidity is added for the first time in the liquidity pool. In these cases, we can confirm that the sniper bots scanned the mempool to swap in the same block where the liquidity is added. However, we also find sniper bots that perform the swap operations a few blocks after the liquidity is created.
@@
We find sniper bots to be less present in Ethereum. Also, in this case, we pick two thresholds and consider all the addresses that swap on average with a distance lower than three blocks (45 seconds) and that swap in at least 10 liquidity pools. We find 64 possible sniper bots that swap in 30% of all the liquidity pools and perform a much smaller fraction of swaps with respect to BSC sniper bots (3.5% of the total). However, interestingly, a higher percentage of swaps are performed in the same block where the liquidity is added in the liquidity pools (60%).
9 1-day Rug Pull Mitigation
Our study highlights that the 1-day rug pulls have some distinctive features. In the following, we propose some metrics that stem from the lessons learned from our analysis that may be useful to build a detection system.
  • Token lifetime: This metric measures the time that elapses since the creation of the token. Indeed, we find that 1-day rug pull operations are performed in a very short timeframe (§7.1).
  • Distribution of the liquidity: This metric tracks the distribution of the LP-tokens. In 1-day rug pulls, the liquidity pool creator owns all the liquidity (§7). Thus, it should be considered extremely risky when a single address owns most of the liquidity.
  • Address rug pull records: This metric tracks addresses that performed a rug pull operation to add them to a list of potential malicious addresses. Indeed, we find that some addresses perform rug pulls multiple times. (§6).
  • Deceptive token name: This metric measures the similarity between the name of tokens contained in the liquidity pools and popular existing tokens or companies. We find that attackers often deceive investors by exploiting the name of the token. (§7.2.3).
An attacker aware of these metrics can try to evade the detection by putting more effort into carrying out the operations (e.g., using different addresses or creating the token in advance). Nonetheless, a distinctive characteristic of 1-day rug pulls is that they are easy to execute and require low effort by the attacker. Thus, we believe the proposed metrics could be sufficient to discourage this operation. Moreover, new metrics and more sophisticated techniques can be developed to identify attackers trying to circumvent the detection. For example, it is possible to follow the money flow between addresses associating different addresses to the same attacker.
We believe that AMMs are interested in leveraging the proposed metrics to build a detection system. Indeed, some have already put effort into this direction. For instance, PancakeSwap recently included in its interface a service called HashDit [33], which provides a risk level in investing in a liquidity pool. HashDit is a Token Contract Scanning service, that estimates the risk of a token by analyzing the code of its smart contract [8]. We believe the proposed metrics can enhance this and other existing services by adding insightful information.
@@ 10 Related Work
Tokens identification. In previous work, there are mainly two token identification approaches: behavior-based and interface-based. The behavior-based method assumes that a token contract maps addresses to the number of tokens owned and contains a function to transfer tokens. Chen et al. [13] follow this approach, analyzing the EVM execution path to find smart contracts data structures that indicate the bookkeeping of a token. The interface-based approach, the technique we take in this work, aims to find tokens that conform to specific interfaces (e.g., the ERC20 interface). This method involves discovering the implemented functions within the smart contract bytecode. Several works use this approach [22, 63, 14]. Frowis et al. [28] proved that the interface-based technique could detect 99% of the tokens in their ground truth dataset.
@@
Liquidity pool scams. Xia et al. [66] characterize scam tokens on Ethereum. First, they leverage CoinMarketCap [18] to obtain a ground truth of official and scam tokens. They used The Graph [31] to obtain 21,778 tokens and 25,131 liquidity pools from May 2020 to December 2020. A guilt-by-association heuristic is adopted to enlarge the dataset, subsequently used to train a machine learning model. More than 11,182 fraudulent tokens were discovered after they ran their classifier on the expanded dataset. Mazorra et al. [43] extended Xia et al. dataset by including Uniswap data until 3 September 2021, discovering an additional 18 thousand scam tokens. They provide three categories for rug pulls: simple, sale, and trap-door. Then, they found that more than 97.7% of the tokens labeled as scams are involved in rug pulls.
Rug pull mitigation. Rug pulls are a very recent issue, and to the best of our knowledge there is no actual solution to prevent them. However, there is a new proposed standard and some protocols that can help to mitigate the problem. To counter the theft of tokens, Wang et al. [64] proposed a new token standard called ERC-20R. With this standard, a transaction is reversible for a short time (dispute period) after it has been performed. During this period, the sender can request to freeze the disputed asset to a set of decentralized judges. If judges agree to lock the disputed asset, it starts another period of time in which the sender can convince judges to revert the transaction. Instead, liquidity locker protocol (e.g., Unicrypt [61]) allows locking LP-tokens inside smart contracts for a given amount of time. This solution assures that the liquidity cannot be removed from the pool until the timer expires, making rug pull impossible. Of course, this solution does not prevent rug pulls after the time expires or dumping one of the tokens in the liquidity pool.
@@ 11 Discussion
What is the impact of not collecting all the internal transactions? Unlike other works [63, 14], we do not collect all smart contracts generated by internal transactions. We collect smart contracts created directly by EOAs, and expand our dataset by adding contracts that emitted at least one Transfer Event. This approach could lead to the loss of a small percentage of tokens. We can perform a rough estimation of the ERC-20 token we miss by comparing the number of tokens we retrieved with the number of tokens retrieved by Chen et al. [14] at the same block height. Our approach retrieves 146,928 tokens instead of 165,955, approximately 12% less. However, it is important to note that, by design, our approach misses only tokens that are never used, traded, or transferred. So, the missing tokens do not represent interesting cases for our study.
@@
Why does it appear that rug pulls and token spammers are more frequent in BSC than in Ethereum? From a technical point of view, rug pulls work the same way in the two blockchains. Indeed, since BSC is EVM compliant and PancakeSwap is a fork of Uniswap, the same smart contract can be used on both blockchains. However, the cost of the operation is significantly different. As we saw in Sec. 7.1.1, performing a rug pull in BSC is cheaper (on average \$10.5 with peaks of \$600) than in Ethereum (on average \$400 with peaks of over \$2,000). These costs represent a fixed cost for the attacker, and going even or gaining money may be more difficult in Ethereum versus BSC.
Are cost-efficient blockchains vulnerable to 1-day rug pulls? As discussed, one of the possible reasons for the prevalence of rug pulls on BSC is the low transaction cost. This could suggest that cost-efficient blockchains are more vulnerable to 1-day rug pulls. However, to confirm this hypothesis, it is necessary to examine whether the phenomenon is common in blockchains with costs similar to the BSC.
Considering our case study of BSC, we believe the low cost of transactions is not the only reason for the high number of rug pulls. In particular, BSC provides one of the first DeFi ecosystems that is cheaper and faster than Ethereum. It quickly became very popular. Moreover, thanks to EVM compatibility, many no-code tools, libraries, and smart contracts already developed for Ethereum can also be used on BSC. This allows the deployment of smart contracts and the creation of new tokens with limited technical capabilities. Thus, the high number of potential victims, the little technical challenge, and the cost-efficiency made the BSC fertile ground for malicious actors to carry out 1-day rug pulls. Even though the low cost can facilitate rug pulls, increasing the costs of blockchains is not a real solution. Instead, a possibility is to shift the focus to DEXes’s protocol and smart contracts for token creation. In particular, it could be possible to design more secure smart contracts to handle tokens (e.g., ERC-20R) or AMM protocols with policies that disincentive rug pull operations.
@@
Can different users coordinate to carry out the same operation, or can a user use multiple addresses? In this work, we considered each address belonging to a single different user, and we assumed there is no coordination among addresses. Nonetheless, a user may change the address he uses to perform each rug pull. It is also possible that a group of users coordinate to carry out the operation. For example, a user can create a liquidity pool while others perform wash trading. A possible approach to detect this malicious behavior is to gather all the transactions among the allegedly involved addresses and look for malicious patterns or communities (e.g., using graph analysis). In this work, we do not perform this analysis, but we plan to explore more sophisticated rug pulls as an extension of this work.